OpenStudy (anonymous):

what is zero of a polynomial function A. the value of the polynoial when zero is substituted for the variable B. the coefficient of the leading term of the polynomial C. a coefficient of the polynomial that is equal to zero D. a value of the variable that makes the polynomial equal to zero

OpenStudy (freckles):

here is an example that might help you to determine your answer for example, let f(x)=x^2-4 sine f(2)=0 then x=2 is a zero because 2^2-4=0 another zero would be x=-2 since f(-2)=0, that is, (-2)^2-4=0 also

OpenStudy (anonymous):

so is it D

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

yeah

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

np also sorry for long wait os is going slow for me :(

OpenStudy (freckles):

I have loading issues anyways